# Inflationary Models with Gauss–Bonnet Coupling in Light of ACT Observations

### v1

@@ 3.3 Hilltop inflation
For hilltop inflation with the potential [66]
$V(\phi)=V_{0}\left[1-\left(\frac{\phi}{\mu}\right)^{p}\right],$ (3.36)
@@
$U(\varphi)=V_{0}\left[1-\left(\frac{\varphi}{\tilde{\mu}}\right)^{p}\right],$ (3.37)
@@
$\epsilon_{U}$ $=\frac{p^{2}\left(\varphi/\tilde{\mu}\right)^{2p-2}}{2\tilde{\mu}^{2}\left[1-\left(\varphi/\tilde{\mu}\right)^{p}\right]^{2}},$ (3.38)
$\eta_{U}$ $=-\frac{(p-1)p\left(\varphi/\tilde{\mu}\right)^{p-2}}{\tilde{\mu}^{2}\left[1-\left(\varphi/\tilde{\mu}\right)^{p}\right]}.$ (3.39)
@@
$n_{s}-1$ $=-\frac{3p^{2}\left(\varphi_{*}/\tilde{\mu}\right)^{2p}}{\varphi_{*}^{2}\left[\left(\varphi_{*}/\tilde{\mu}\right)^{p}-1\right]^{2}}+\frac{2(p-1)p\left(\varphi_{*}/\tilde{\mu}\right)^{p}}{\varphi_{*}^{2}\left[\left(\varphi_{*}/\tilde{\mu}\right)^{p}-1\right]},$ (3.42)
$r$ $=\frac{16(1-\lambda)p^{2}\left(\varphi_{*}/\tilde{\mu}\right)^{2p-2}}{2\tilde{\mu}^{2}\left[1-\left(\varphi_{*}/\tilde{\mu}\right)^{p}\right]^{2}}.$ (3.43)
@@
$n_{s}=1-\frac{2(p-1)}{(p-2)N},\quad r=\frac{8p^{2}}{\tilde{\mu}^{2}}\left[\frac{\tilde{\mu}^{2}}{p(p-2)N}\right]^{(2p-2)/(p-2)}.$ (3.44)
@@
By numerically solving the scalar spectral index for a general $\tilde{\mu}$, and comparing it with the observational constraints given in Eq. (3.14), we obtain the constraints on the parameters of the hilltop inflation model, shown in Fig. 4. For the case $p=4$, the predictions of the hilltop inflation model with varying $\tilde{\mu}$ are presented in Fig. 1, represented by the green curves. With the inclusion of the Gauss-Bonnet term and a coupling constant $\lambda=0.8$, the model can be brought into agreement with the P-ACT-LB-BK18 observational data for sufficiently large values of $\tilde{\mu}$.
Figure 4: Constraints on the parameters $p$ and $\tilde{\mu}$ of the hilltop inflation model, derived from the P-ACT-LB data as given in Eq. (3.14)
4 Conclusion
@@
Our results suggest that the Gauss–Bonnet coupling may help bring a broad class of inflationary models into better agreement with current CMB measurements. This mechanism offers a potentially useful approach for easing tensions between theoretical predictions and observational data, and its relevance is likely to extend beyond the specific examples considered here.
Acknowledgments
This work is supported in part by the National Natural Science Foundation of China under Grant No. 12205015, and the National Key Research and Development Program of China under Grant No. 2020YFC2201504.

### v2

@@ 3.3 Hilltop inflation
For hilltop inflation with the potential Boubekeur:2005zm
$V(\phi)=V_{0}\left[1-\left(\frac{\phi}{\mu}\right)^{p}\right],$ (68)
@@
$U(\varphi)=V_{0}\left[1-\left(\frac{\varphi}{\tilde{\mu}}\right)^{p}\right],$ (69)
@@
$\epsilon_{U}$ $=\frac{p^{2}\left(\varphi/\tilde{\mu}\right)^{2p-2}}{2\tilde{\mu}^{2}\left[1-\left(\varphi/\tilde{\mu}\right)^{p}\right]^{2}},$ (70)
$\eta_{U}$ $=-\frac{(p-1)p\left(\varphi/\tilde{\mu}\right)^{p-2}}{\tilde{\mu}^{2}\left[1-\left(\varphi/\tilde{\mu}\right)^{p}\right]}.$ (71)
@@
$n_{s}$ $=1-\frac{3p^{2}\left(\varphi_{*}/\tilde{\mu}\right)^{2p}}{\varphi_{*}^{2}\left[\left(\varphi_{*}/\tilde{\mu}\right)^{p}-1\right]^{2}}+\frac{2(p-1)p\left(\varphi_{*}/\tilde{\mu}\right)^{p}}{\varphi_{*}^{2}\left[\left(\varphi_{*}/\tilde{\mu}\right)^{p}-1\right]},$ (74)
$r$ $=\frac{16(1-\lambda)p^{2}\left(\varphi_{*}/\tilde{\mu}\right)^{2p-2}}{2\tilde{\mu}^{2}\left[1-\left(\varphi_{*}/\tilde{\mu}\right)^{p}\right]^{2}}.$ (75)
@@
$n_{s}=1-\frac{2(p-1)}{(p-2)N},$ (76)
$r=\frac{8p^{2}}{\tilde{\mu}^{2}}\left[\frac{\tilde{\mu}^{2}}{p(p-2)N}\right]^{(2p-2)/(p-2)}.$ (77)
@@
By numerically solving the scalar spectral index for a general $\tilde{\mu}$, and comparing it with the observational constraints given in Eq. (46), we obtain the constraints on the parameters of the hilltop inflation model, shown in Fig. 5. For the case $p=4$, the predictions of the hilltop inflation model with varying $\tilde{\mu}$ are presented in Fig. 2, represented by the green curves. With the inclusion of the Gauss-Bonnet term and a coupling constant $\lambda=0.8$, the model can be brought into agreement with the P-ACT-LB-BK18 observational data for sufficiently large values of $\tilde{\mu}$.
Figure 5: Constraints on the parameters $p$ and $\tilde{\mu}$ of the hilltop inflation model, derived from the P-ACT-LB data as given in Eq. (46). The blue region corresponds to parameter values for which the predicted scalar spectral index given by Eq. (74) from the hilltop inflation is consistent with the P-ACT-LB data.
Similar results can also be obtained for other values in the moderate range $0.5<\lambda<1$, which indicates that the agreement with the observational constraints does not rely on a fine-tuned choice of $\lambda$.
4 Discuss and Conclusion
@@
Our results suggest that the Gauss–Bonnet coupling may help bring a broad class of inflationary models into better agreement with current CMB measurements. This mechanism offers a potentially useful approach for easing tensions between theoretical predictions and observational data, and its relevance is likely to extend beyond the specific examples considered here. We also note that, although the suppression of the tensor-to-scalar ratio becomes more effective as $\lambda\to 1$, our mechanism does not rely on extreme fine-tuning: values in the moderate range $0.5<\lambda<1$ are already sufficient to satisfy the current observational limits. Nevertheless, if $|1-\lambda|$ becomes smaller than the first-order slow-roll parameters, higher-order corrections should be considered beyond the approximation used in this work.
Upcoming CMB polarization experiments such as Lite (Light) satellite for the studies of B-mode polarization and Inflation from cosmic background Radiation Detection (LiteBIRD) LiteBIRD:2022cnt and CMB-S4 CMB-S4:2016ple will dramatically improve the sensitivity to primordial tensor modes, aiming at the level of $r\lesssim 10^{-3}$. If primordial gravitational waves are detected and the tensor-to-scalar ratio is found in the range $10^{-2}$$10^{-3}$, such an observation would support the Gauss–Bonnet suppression mechanism, corresponding to $1-\lambda\sim\mathcal{O}(10^{-1})$. While this may raise concerns about fine-tuning, we note that $1-\lambda$ at this level still exceeds the first-order slow-roll parameters, so that our analytic results remain valid within the slow-roll framework. Conversely, a null result at $r<10^{-3}$ would push the model into a regime with $1-\lambda<\mathcal{O}(10^{-2})$, where higher-order corrections may become necessary. In addition, future CMB observations will further improve the precision of the scalar spectral index $n_{s}$, which may provide complementary constraints on Gauss–Bonnet inflationary scenarios.
Acknowledgements.
This work is supported in part by the National Key Research and Development Program of China under Grant No. 2020YFC2201504, and the National Natural Science Foundation of China under Grant No. 12205015.